# SuperQuadricOcc: Multi-Layer Gaussian Approximation of Superquadrics for Real-Time Self-Supervised Occupancy Estimation

### v1

@@ Abstract
Semantic occupancy estimation enables comprehensive scene understanding for automated driving, providing dense spatial and semantic information essential for perception and planning. While Gaussian representations have been widely adopted in self-supervised occupancy estimation, the deployment of a large number of Gaussian primitives drastically increases memory requirements and is not suitable for real-time inference. In contrast, superquadrics permit reduced primitive count and lower memory requirements due to their diverse shape set. However, implementation into a self-supervised occupancy model is nontrivial due to the absence of a superquadric rasterizer to enable model supervision. Our proposed method, SuperQuadricOcc, employs a superquadric-based scene representation. By leveraging a multi-layer icosphere-tessellated Gaussian approximation of superquadrics, we enable Gaussian rasterization for supervision during training. On the Occ3D dataset, SuperQuadricOcc achieves a 75% reduction in memory footprint, 124% faster inference, and a 5.9% improvement in mIoU compared to previous Gaussian-based methods, without the use of temporal labels. To our knowledge, this is the first occupancy model to enable real-time inference while maintaining competitive performance. The use of superquadrics reduces the number of primitives required for scene modeling by 84% relative to Gaussian-based approaches. Finally, evaluation against prior methods is facilitated by our fast superquadric voxelization module. The code will be released as open source.
@@ I Introduction
Figure 1: Comparison of GaussianFlowOcc [boeder2025gaussianflowocc] and our model, SuperQuadricOcc: Replacing Gaussians with superquadrics enables more efficient scene modeling through a sparser and more expressive representation. Our method achieves higher mIoU, faster inference, and lower memory consumption with significantly fewer primitives. Ground truth is provided in Figure 4.
@@
The application of computer vision to scene understanding has gained significant traction, particularly in perception for automated vehicles, which focuses on estimating dynamic and complex driving environments. Initially, 3D object detection served as the foundation for spatial understanding, providing precise localization of agents in the scene [mao2023_3d_bounding_survey]. However, its cuboid prediction space limits the representation of irregular objects, such as bicycles. Bird’s Eye View (BEV) later emerged as a top-down fixed-grid prediction approach offering broader scene coverage [ma2024_vision_bev_survey], but its lack of vertical reasoning restricted comprehensive 3D understanding, thus limiting real-world applicability. Semantic occupancy estimation addressed this by extending BEV into 3D, representing the world as a discretized voxel grid.
@@
Within semantic occupancy estimation, scene representation has evolved from dense voxel-based formats to more efficient Gaussian representations, and most recently, to superquadrics [zuo2025quadricformer]; however, superquadrics have not yet been explored in the self-supervised setting. Self-supervised occupancy models rely on 2D pseudo-labels derived from Visual Foundation Models (VFMs), such as Grounded-SAM [ren2024grounded, kirillov2023segment, liu2024grounding] for semantic information and Metric3Dv2 [hu2024metric3d] for depth cues, eliminating the need for dense manual annotations [jiang2025gausstr, gaussianocc, zhang2023occnerf]. Given the discrepancy between the 2D supervision space and the 3D evaluation space, neural rendering methods such as NeRF [mildenhall2021nerf] and Gaussian Splatting [kerbl20233dgs] are well-suited for bridging this gap. Nonetheless, challenges remain, including reliance on large foundation models such as Metric3Dv2 during inference to address depth ambiguity [jiang2025gausstr], and the need for thousands of Gaussian primitives to model fine details [boeder2025gaussianflowocc]. Both factors increase inference time and memory requirements, hindering real-time deployment.
@@
In an effort to achieve real-time inference, we introduce the self-supervised occupancy model SuperQuadricOcc. Our approach employs superquadrics for scene representation, leveraging their diverse shape set to model objects more compactly than previous Gaussian-based methods. This design reduces the number of primitives required for scene modeling by 84% compared to GaussianFlowOcc [boeder2025gaussianflowocc]. For efficient and effective supervision against 2D labels, we apply a multi-layer Gaussian approximation to each superquadric primitive, enabling Gaussian rasterization as a proxy for rendering the superquadric distribution.
@@
Incorporating temporal labels for adjacent renderings, as done in GaussianFlowOcc [boeder2025gaussianflowocc], improves evaluation metrics by enhancing view consistency. However, we omit this component in our model and in GaussianFlowOcc for comparison, as it more than triples the training time. As illustrated in Figure 1, superquadrics model the scene more sparsely yet provide a richer and more structured representation compared to Gaussians. This translates directly into improved efficiency and accuracy. SuperQuadricOcc achieves a 12.69 mean Intersection over Union (mIoU) on the Occ3D-nuScenes dataset [tian2023occ3d], marking a 5.9% improvement over GaussianFlowOcc. Additionally, due to the compact scene representation, our model attains real-time inference at 21.5 FPS. Memory usage is also significantly reduced, requiring only 0.70 GB during inference, a 75% reduction. By exploiting the efficiency of superquadrics, our model delivers real-time inference with strong performance, establishing a foundation for future research in superquadric-based occupancy estimation.
@@ II-A Semantic Occupancy Prediction
Semantic occupancy estimation can generally be categorized into two main approaches: supervised and self-supervised. Both approaches aim to estimate voxel-level semantics in a discretized 3D environment. Supervised occupancy estimation has seen significant growth with the introduction of standardized datasets [wei2023surroundocc, tian2023occ3d, wang2023openoccupancy, sima2023scene, li2024sscbench], which provide occupancy labels for major benchmark datasets such as Waymo [sun2020waymo], nuScenes [nuScenes], and KITTI [liao2022kitti]. Scene representation relies on voxel-based structure largely due to the voxel-based evaluation space [li2024bevformer, renderworld, ming2024occfusion, wang2024panoocc, wu2025deep]. However, alternative representations have emerged, including cylindrical-coordinate formats [ming2025occcylindrical], octree-based structures [lu2024octreeocc], and, more recently, Gaussian representations [huang2024gaussianformer, huang2025gaussianformer2, zuo2025gaussianworld, zhao2025gaussianformer3d]. Gaussian scene representation has become particularly popular due to its reduced memory requirements compared to dense voxel grids [huang2024gaussianformer], and its compatibility with Gaussian rasterization, which enables additional loss computation [park2025s2go, shi2025odg] for view consistency, improving overall scene understanding. Furthermore, Gaussian rasterization has been integrated into voxel-based models as a supplementary loss through voxel-to-Gaussian conversion modules [renderworld, gaussianpretrain, sun2024gsrender, lei2025tacocc, chambon2025gaussrender].
In self-supervised models, Gaussian scene representations are particularly advantageous due to this research domain’s reliance on 2D labels for supervision [gaussianocc, jiang2025gausstr, boeder2025gaussianflowocc]. GaussianOcc first demonstrated this by extending OccNeRF [zhang2023occnerf], replacing the NeRF renderer with Gaussian splatting; however, both approaches suffer from high memory consumption due to their voxel-based structures. GaussTR [jiang2025gausstr] introduced a foundation model–style framework, incorporating VFMs such as Talk2DINO [barsellotti2024talking], CLIP [radford2021learning], Metric3Dv2 [hu2024metric3d], and DINOv2 [oquab2023dinov2], achieving state-of-the-art performance. Nonetheless, its reliance on Metric3Dv2 during inference and inefficient voxelization techniques limits real-time applicability. GaussianFlowOcc [boeder2025gaussianflowocc] further advanced prior work by introducing a temporal flow module for improved view consistency and induced attention for faster and more effective interactions between Gaussian features. However, the use of 10,000 Gaussian primitives substantially increases inference time, again preventing real-time performance. Furthermore, incorporating temporal labels via the temporal flow module triples training time and is therefore omitted from our experiments.
@@ II-B SuperQuadrics for Scene Modelling
Superquadrics were first introduced and parameterized by Alan H. Barr [barr1981superquadrics] as a unified mathematical framework for applications in computer graphics and numerical methods for 3D design. In computer vision, superquadrics have been widely adopted for scene modeling, particularly in 3D reconstruction [alaniz2023iterative, zhang2003superquadric, paschalidou2019superquadrics, gao2024partgs, jiang2024gaussianblock] and static scene reconstruction [fedele2025superdec], where neural rendering techniques are often used to provide supervision.
Paschalidou et al. [paschalidou2019superquadrics] proposed representing 3D shapes as compositions of superquadrics rather than cuboids. The model is trained in an unsupervised manner using a bidirectional Chamfer distance between points sampled from predicted superquadric surfaces and ground-truth occupancy reconstructions. PartGS [gao2024partgs] and GaussianBlock [jiang2024gaussianblock] further extended this approach, drawing inspiration from prior work that places Gaussians on mesh surfaces for rendering [guedon2024sugar, waczynska2024games]. Both methods approximate the superquadric surface by positioning Gaussians on a deformed icosphere whose vertices lie on the surface of the superquadric. In PartGS, both the Gaussians and superquadric geometry are jointly optimized, while in GaussianBlock, the superquadric primitives are first learned via differentiable rendering and later enhanced with bound Gaussians for high-fidelity rendering and editing. SuperDec [fedele2025superdec] fits superquadrics to point cloud data using a Point-Voxel CNN and a transformer decoder to predict superquadric parameters.
The use of superquadrics in supervised occupancy estimation was first explored in QuadricFormer [zuo2025quadricformer], which outperforms the Gaussian-based GaussianFormer [huang2024gaussianformer] in evaluation metrics while reducing inference time and memory consumption. In our model, SuperQuadricOcc, we extend this concept to self-supervised occupancy estimation, marking the first application of superquadrics in this setting. For model supervision, we adopt the icosphere tessellation strategy from PartGS [gao2024partgs] and GaussianBlock [jiang2024gaussianblock], approximating the superquadric surface with Gaussians. Since we model superquadrics as occupancy probabilities, we further this approach by introducing a multi-layer Gaussian formulation to effectively capture the underlying distribution.
Figure 2: Architecture of SuperQuadricOcc: The model extracts image features and refines learnable superquadric features and means in the Transformer Network, similar to GaussianFlowOcc [boeder2025gaussianflowocc]. The MLP prediction heads estimate the remaining superquadric properties: $\{\bm{s},\sigma,\bm{c},\varepsilon_{1},\varepsilon_{2}\}$. During training, the SuperQuadric to Gaussian Module generates a multi-layer Gaussian approximation per superquadric for rendered supervision, while during inference, the SuperQuadric Voxelizer efficiently produces voxel labels for evaluation.
@@ III-B SuperQuadrics
As described by Barr [barr1981superquadrics], the $\varepsilon$ parameters control the squareness and roundness of the shape: $\varepsilon_{1}$ governs squareness along the north–south direction, while $\varepsilon_{2}$ does so along the east–west direction. The inside–outside function, $f$, for a superquadric is defined as:
@@
where $\mathbf{x}=(x,y,z)^{T}$ and $\mathbf{s}=(s_{x},s_{y},s_{z})^{T}$. Following Zuo et al. [zuo2025quadricformer], the occupancy probability of a superquadric can be formulated as follows. To evaluate a point $\mathbf{x}$ for a given superquadric $S$, it is first transformed into the local coordinate system of the superquadric via rotation and translation to obtain $\mathbf{x}_{\mathbf{S}}$:
@@ III-D SuperQuadric to Gaussian Module
To enable rendering of superquadrics, we adopt a multi-layer Gaussian approximation. Previous works approximated only the surface of superquadrics using 2D Gaussians for 3D modeling [gao2024partgs, jiang2024gaussianblock]. In contrast, since our formulation treats superquadrics as occupancy probabilities, we extend this approach with a multi-layer approximation using 3D Gaussians.
@@
We approximate each occupancy distribution of the scaled superquadrics following established frameworks for Gaussian placement on mesh surfaces [gao2024partgs, jiang2024gaussianblock]. For each $S_{k}\in S_{K}$, a tessellated unit icosphere is constructed, with vertices lying on the surface of a unit sphere. Each vertex on the icosphere, defined by spherical coordinates $(\eta,\omega)$, is mapped to the superquadric surface through signed powers of trigonometric functions, controlled by $\varepsilon_{1}$ and $\varepsilon_{2}$. This produces a mesh within the superquadric that approximates its surface. Icospheres with a higher vertex count yield a more accurate approximation of the superquadric surface at the cost of additional computation. Gaussians are then centered on each face center and oriented along the surface normal and tangent directions, with their $xy$ scales proportional to the local face area. This is performed so that each Gaussian sufficiently covers the area of each face. Each Gaussian inherits the semantics of its parent superquadric. For Gaussian opacity, we deviate from prior formulations, which assign the opacity of each Gaussian to that of the parent superquadric, and we instead assign it to be:
@@ III-E SuperQuadric Voxelization
To enable direct comparison with previous state-of-the-art methods, we develop a superquadric voxelization module for evaluation. Following the procedure of Boeder et al. [boeder2025gaussianflowocc], we voxelize opacity and semantics separately, to obtain the following formulation:
@@
For each voxel $v$ with center $\bm{p}$ in the discretized grid, we aggregate contributions from all superquadrics $S$ within its neighborhood $N=5$, similar to GaussianFlowOcc [boeder2025gaussianflowocc]. Each superquadric is then transformed into the local coordinate frame of the voxel using Equation (3) to obtain $\bm{p_{s}}$. The contribution is computed using the weighted sums of the occupancy probability function defined in Equation (4). Voxels below a defined opacity threshold $\tau$, are assigned as free in the semantic voxel to obtain the final prediction. We assign $\tau=0.01$, through empirical tuning.
TABLE I: Occupancy Estimation results on the Occ3D-nuScenes [tian2023occ3d] dataset: Rep. denotes the method of scene representation: Gauss - Gaussian; Quad - SuperQuadric. Mem. denotes the memory (GB) required for a single sample inference. GaussianFlowOcc results are reproduced locally without the deployment of temporal labels. GaussianFlowOcc* denotes the equivalent trained model with 1,600 Gaussians. The best-performing results between the GaussianFlowOcc models and SuperQuadricOcc only are highlighted in bold. Other state-of-the-art models are included for reference.
Model Rep. Mem. FPS IoU mIoU $\blacksquare$ barrier $\blacksquare$ bicycle $\blacksquare$ bus $\blacksquare$ car $\blacksquare$ const. veh. $\blacksquare$ motorcycle $\blacksquare$ pedestrian $\blacksquare$ traffic cone $\blacksquare$ trailer $\blacksquare$ truck $\blacksquare$ drive. surf. $\blacksquare$ sidewalk $\blacksquare$ terrain $\blacksquare$ manmade $\blacksquare$ vegetation SelfOcc [huang2024selfocc] Voxel 3.00 7.4 45.01 10.54 0.15 0.66 5.46 12.54 0.00 0.80 2.10 0.00 0.00 8.25 55.49 26.30 26.54 14.22 5.60 OccNeRF [zhang2023occnerf] Voxel 12.68 5.4 46.43 10.81 0.83 0.82 5.13 12.49 3.50 0.23 3.10 1.84 0.52 3.90 52.62 20.81 24.75 18.45 13.19 GaussianOcc [gaussianocc] Voxel 11.81 5.4 42.91 11.26 1.79 5.82 14.58 13.55 1.30 2.82 7.95 9.76 0.56 9.61 44.59 20.10 17.58 8.61 10.29 GaussTR [jiang2025gausstr] Gauss 1.97 0.3 44.54 13.76 6.50 8.54 21.77 24.27 6.26 15.48 7.94 1.86 6.10 17.16 36.98 17.21 7.16 21.18 9.99 GaussianFlowOcc [boeder2025gaussianflowocc] Gauss 2.85 9.6 35.85 11.98 5.48 1.85 9.26 9.41 0.33 1.88 3.00 3.03 1.51 7.32 57.77 22.26 28.25 13.59 14.72 GaussianFlowOcc* [boeder2025gaussianflowocc] Gauss 0.62 20.4 35.40 9.98 3.20 1.60 8.06 7.40 1.18 2.51 2.14 1.93 1.77 6.09 47.34 19.90 22.52 13.22 10.88 SuperQuadricOcc (Ours) Quad 0.70 21.5 33.67 12.69 5.78 5.69 16.32 13.06 3.72 5.69 3.94 3.54 2.56 11.94 43.58 20.66 24.15 14.45 15.29

### v3

@@ Abstract
Semantic occupancy estimation enables comprehensive scene understanding for automated driving, providing dense spatial and semantic information essential for perception and planning. While Gaussian representations have been widely adopted in self-supervised occupancy estimation, the deployment of a large number of Gaussian primitives drastically increases memory requirements and is not suitable for real-time inference. In contrast, superquadrics permit reduced primitive count and lower memory requirements due to their diverse shape set. However, implementation into a self-supervised occupancy model is nontrivial due to the absence of a superquadric rasterizer to enable model supervision. Our proposed method, SuperQuadricOcc, employs a superquadric-based scene representation. By leveraging a multi-layer icosphere-tessellated Gaussian approximation of superquadrics, we enable Gaussian rasterization for supervision during training. On the Occ3D dataset, SuperQuadricOcc achieves a 75% reduction in memory footprint, 124% faster inference, and a 5.9% improvement in mIoU compared to previous Gaussian-based methods, without the use of temporal labels. To our knowledge, this is the first occupancy model to enable real-time inference while maintaining competitive performance. The use of superquadrics reduces the number of primitives required for scene modeling by 84% relative to Gaussian-based approaches. Finally, evaluation against prior methods is facilitated by our fast superquadric voxelization module. The code will be made available at https://github.com/seamie6/SuperQuadricOcc.
@@ I Introduction
Figure 1: Comparison of GaussianFlowOcc [4] and our model, SuperQuadricOcc: Replacing Gaussians with superquadrics enables more efficient scene modeling through a sparser and more expressive representation. Our method achieves higher mIoU, faster inference, and lower memory consumption with significantly fewer primitives. Ground truth is provided in Figure 4.
@@
The application of computer vision to scene understanding has gained significant traction, particularly in perception for automated vehicles, which focuses on estimating dynamic and complex driving environments. Initially, 3D object detection served as the foundation for spatial understanding, providing precise localization of agents in the scene [26]. However, its cuboid prediction space limits the representation of irregular objects, such as bicycles. Bird’s Eye View (BEV) later emerged as a top-down fixed-grid prediction approach offering broader scene coverage [25], but its lack of vertical reasoning restricted comprehensive 3D understanding, thus limiting real-world applicability. Semantic occupancy estimation addressed this by extending BEV into 3D, representing the world as a discretized voxel grid.
@@
Within semantic occupancy estimation, scene representation has evolved from dense voxel-based formats to more efficient Gaussian representations, and most recently, to superquadrics [51]; however, superquadrics have not yet been explored in the self-supervised setting. Self-supervised occupancy models rely on 2D pseudo-labels derived from Visual Foundation Models (VFMs), such as Grounded-SAM [34, 18, 23] for semantic information and Metric3Dv2 [11] for depth cues, eliminating the need for dense manual annotations [15, 8, 48]. Given the discrepancy between the 2D supervision space and the 3D evaluation space, neural rendering methods such as NeRF [27] and Gaussian Splatting [17] are well-suited for bridging this gap. Nonetheless, challenges remain, including reliance on large foundation models such as Metric3Dv2 during inference to address depth ambiguity [15], and the need for thousands of Gaussian primitives to model fine details [4]. Both factors increase inference time and memory requirements, hindering real-time deployment.
@@
In an effort to achieve real-time inference, we introduce the self-supervised occupancy model SuperQuadricOcc. Our approach employs superquadrics for scene representation, leveraging their diverse shape set to model objects more compactly than previous Gaussian-based methods. This design reduces the number of primitives required for scene modeling by 84% compared to GaussianFlowOcc [4]. For efficient and effective supervision against 2D labels, we apply a multi-layer Gaussian approximation to each superquadric primitive, enabling Gaussian rasterization as a proxy for rendering the superquadric distribution.
@@
Incorporating temporal labels for adjacent renderings, as done in GaussianFlowOcc [4], improves evaluation metrics by enhancing view consistency. However, we omit this component in our model and in GaussianFlowOcc for comparison, as it more than triples the training time. As illustrated in Figure 1, superquadrics model the scene more sparsely yet provide a richer and more structured representation compared to Gaussians. This translates directly into improved efficiency and accuracy. SuperQuadricOcc achieves a 12.69 mean Intersection over Union (mIoU) on the Occ3D-nuScenes dataset [40], marking a 5.9% improvement over GaussianFlowOcc. Additionally, due to the compact scene representation, our model attains real-time inference at 21.5 FPS. Memory usage is also significantly reduced, requiring only 0.70 GB during inference, a 75% reduction. By exploiting the efficiency of superquadrics, our model delivers real-time inference with strong performance, establishing a foundation for future research in superquadric-based occupancy estimation.
@@ II-A Semantic Occupancy Prediction
Semantic occupancy estimation can generally be categorized into two main approaches: supervised and self-supervised. Both approaches aim to estimate voxel-level semantics in a discretized 3D environment. Supervised occupancy estimation has seen significant growth with the introduction of standardized datasets [44, 40, 42, 36, 20], which provide occupancy labels for major benchmark datasets such as Waymo [37], nuScenes [5], and KITTI [22]. Scene representation relies on voxel-based structure largely due to the voxel-based evaluation space [21, 47, 29, 43, 45]. However, alternative representations have emerged, including cylindrical-coordinate formats [28], octree-based structures [24], and, more recently, Gaussian representations [14, 12, 52, 50]. Gaussian scene representation has become particularly popular due to its reduced memory requirements compared to dense voxel grids [14], and its compatibility with Gaussian rasterization, which enables additional loss computation [31, 35] for view consistency, improving overall scene understanding. Furthermore, Gaussian rasterization has been integrated into voxel-based models as a supplementary loss through voxel-to-Gaussian conversion modules [47, 46, 38, 19, 6].
In self-supervised models, Gaussian scene representations are particularly advantageous due to this research domain’s reliance on 2D labels for supervision [8, 15, 4]. GaussianOcc first demonstrated this by extending OccNeRF [48], replacing the NeRF renderer with Gaussian splatting; however, both approaches suffer from high memory consumption due to their voxel-based structures. GaussTR [15] introduced a foundation model–style framework, incorporating VFMs such as Talk2DINO [3], CLIP [33], Metric3Dv2 [11], and DINOv2 [30], achieving state-of-the-art performance. Nonetheless, its reliance on Metric3Dv2 during inference and inefficient voxelization techniques limits real-time applicability. GaussianFlowOcc [4] further advanced prior work by introducing a temporal flow module for improved view consistency and induced attention for faster and more effective interactions between Gaussian features. However, the use of 10,000 Gaussian primitives substantially increases inference time, again preventing real-time performance. Furthermore, incorporating temporal labels via the temporal flow module triples training time and is therefore omitted from our experiments.
@@ II-B SuperQuadrics for Scene Modelling
Superquadrics were first introduced and parameterized by Alan H. Barr [2] as a unified mathematical framework for applications in computer graphics and numerical methods for 3D design. In computer vision, superquadrics have been widely adopted for scene modeling, particularly in 3D reconstruction [1, 49, 32, 9, 16] and static scene reconstruction [7], where neural rendering techniques are often used to provide supervision.
Paschalidou et al. [32] proposed representing 3D shapes as compositions of superquadrics rather than cuboids. The model is trained in an unsupervised manner using a bidirectional Chamfer distance between points sampled from predicted superquadric surfaces and ground-truth occupancy reconstructions. PartGS [9] and GaussianBlock [16] further extended this approach, drawing inspiration from prior work that places Gaussians on mesh surfaces for rendering [10, 41]. Both methods approximate the superquadric surface by positioning Gaussians on a deformed icosphere whose vertices lie on the surface of the superquadric. In PartGS, both the Gaussians and superquadric geometry are jointly optimized, while in GaussianBlock, the superquadric primitives are first learned via differentiable rendering and later enhanced with bound Gaussians for high-fidelity rendering and editing. SuperDec [7] fits superquadrics to point cloud data using a Point-Voxel CNN and a transformer decoder to predict superquadric parameters.
The use of superquadrics in supervised occupancy estimation was first explored in QuadricFormer [51], which outperforms the Gaussian-based GaussianFormer [14] in evaluation metrics while reducing inference time and memory consumption. In our model, SuperQuadricOcc, we extend this concept to self-supervised occupancy estimation, marking the first application of superquadrics in this setting. For model supervision, we adopt the icosphere tessellation strategy from PartGS [9] and GaussianBlock [16], approximating the superquadric surface with Gaussians. Since we model superquadrics as occupancy probabilities, we further this approach by introducing a multi-layer Gaussian formulation to effectively capture the underlying distribution.
Figure 2: Architecture of SuperQuadricOcc: The model extracts image features and refines learnable superquadric features and means in the Transformer Network, similar to GaussianFlowOcc [4]. The MLP prediction heads estimate the remaining superquadric properties: $\{\bm{s},\sigma,\bm{c},\varepsilon_{1},\varepsilon_{2}\}$. During training, the SuperQuadric to Gaussian Module generates a multi-layer Gaussian approximation per superquadric for rendered supervision, while during inference, the SuperQuadric Voxelizer efficiently produces voxel labels for evaluation.
@@ III-B SuperQuadrics
As described by Barr [2], the $\varepsilon$ parameters control the squareness and roundness of the shape: $\varepsilon_{1}$ governs squareness along the north–south direction, while $\varepsilon_{2}$ does so along the east–west direction. The inside–outside function, $f$, for a superquadric is defined as:
@@
where $\mathbf{x}=(x,y,z)^{T}$ and $\mathbf{s}=(s_{x},s_{y},s_{z})^{T}$. Following Zuo et al. [51], the occupancy probability of a superquadric can be formulated as follows. To evaluate a point $\mathbf{x}$ for a given superquadric $S$, it is first transformed into the local coordinate system of the superquadric via rotation and translation to obtain $\mathbf{x}_{\mathbf{S}}$:
@@ III-D SuperQuadric to Gaussian Module
To enable rendering of superquadrics, we adopt a multi-layer Gaussian approximation. Previous works approximated only the surface of superquadrics using 2D Gaussians for 3D modeling [9, 16]. In contrast, since our formulation treats superquadrics as occupancy probabilities, we extend this approach with a multi-layer approximation using 3D Gaussians.
@@
We approximate each occupancy distribution of the scaled superquadrics following established frameworks for Gaussian placement on mesh surfaces [9, 16]. For each $S_{k}\in S_{K}$, a tessellated unit icosphere is constructed, with vertices lying on the surface of a unit sphere. Each vertex on the icosphere, defined by spherical coordinates $(\eta,\omega)$, is mapped to the superquadric surface through signed powers of trigonometric functions, controlled by $\varepsilon_{1}$ and $\varepsilon_{2}$. This produces a mesh within the superquadric that approximates its surface. Icospheres with a higher vertex count yield a more accurate approximation of the superquadric surface at the cost of additional computation. Gaussians are then centered on each face center and oriented along the surface normal and tangent directions, with their $xy$ scales proportional to the local face area. This is performed so that each Gaussian sufficiently covers the area of each face. Each Gaussian inherits the semantics of its parent superquadric. For Gaussian opacity, we deviate from prior formulations, which assign the opacity of each Gaussian to that of the parent superquadric, and we instead assign it to be:
@@ III-E SuperQuadric Voxelization
To enable direct comparison with previous state-of-the-art methods, we develop a superquadric voxelization module for evaluation. Following the procedure of Boeder et al. [4], we voxelize opacity and semantics separately, to obtain the following formulation:
@@
For each voxel $v$ with center $\bm{p}$ in the discretized grid, we aggregate contributions from all superquadrics $S$ within its neighborhood $N=5$, similar to GaussianFlowOcc [4]. Each superquadric is then transformed into the local coordinate frame of the voxel using Equation (3) to obtain $\bm{p_{s}}$. The contribution is computed using the weighted sums of the occupancy probability function defined in Equation (4). Voxels below a defined opacity threshold $\tau$, are assigned as free in the semantic voxel to obtain the final prediction. We assign $\tau=0.01$, through empirical tuning.
TABLE I: Occupancy Estimation results on the Occ3D-nuScenes [40] dataset: Rep. denotes the method of scene representation: Gauss - Gaussian; Quad - SuperQuadric. Mem. denotes the memory (GB) required for a single sample inference. GaussianFlowOcc results are reproduced locally without the deployment of temporal labels. GaussianFlowOcc* denotes the equivalent trained model with 1,600 Gaussians. The best-performing results between the GaussianFlowOcc models and SuperQuadricOcc only are highlighted in bold. Other state-of-the-art models are included for reference.
Model Rep. Mem. FPS IoU mIoU $\blacksquare$ barrier $\blacksquare$ bicycle $\blacksquare$ bus $\blacksquare$ car $\blacksquare$ const. veh. $\blacksquare$ motorcycle $\blacksquare$ pedestrian $\blacksquare$ traffic cone $\blacksquare$ trailer $\blacksquare$ truck $\blacksquare$ drive. surf. $\blacksquare$ sidewalk $\blacksquare$ terrain $\blacksquare$ manmade $\blacksquare$ vegetation SelfOcc [13] Voxel 3.00 7.4 45.01 10.54 0.15 0.66 5.46 12.54 0.00 0.80 2.10 0.00 0.00 8.25 55.49 26.30 26.54 14.22 5.60 OccNeRF [48] Voxel 12.68 5.4 46.43 10.81 0.83 0.82 5.13 12.49 3.50 0.23 3.10 1.84 0.52 3.90 52.62 20.81 24.75 18.45 13.19 GaussianOcc [8] Voxel 11.81 5.4 42.91 11.26 1.79 5.82 14.58 13.55 1.30 2.82 7.95 9.76 0.56 9.61 44.59 20.10 17.58 8.61 10.29 GaussTR [15] Gauss 1.97 0.3 44.54 13.76 6.50 8.54 21.77 24.27 6.26 15.48 7.94 1.86 6.10 17.16 36.98 17.21 7.16 21.18 9.99 GaussianFlowOcc [4] Gauss 2.85 9.6 35.85 11.98 5.48 1.85 9.26 9.41 0.33 1.88 3.00 3.03 1.51 7.32 57.77 22.26 28.25 13.59 14.72 GaussianFlowOcc* [4] Gauss 0.62 20.4 35.40 9.98 3.20 1.60 8.06 7.40 1.18 2.51 2.14 1.93 1.77 6.09 47.34 19.90 22.52 13.22 10.88 SuperQuadricOcc (Ours) Quad 0.70 21.5 33.67 12.69 5.78 5.69 16.32 13.06 3.72 5.69 3.94 3.54 2.56 11.94 43.58 20.66 24.15 14.45 15.29